OpenStudy (anonymous):

Given the points (-2, 1) and (5, -5): a) Find the slope of the line through the points. b) Write an equation in point-slope form of the line through the points. c) Convert the equation to slope-intercept form. d) Convert the equation to standard form. e) Graph the equation. You may use the axes provided, or create your own graph.

OpenStudy (anonymous):

R These Yo Answers Or What Youu Suppose To Do ,

OpenStudy (anonymous):

?

OpenStudy (anonymous):

these arent the answers the question is Given the points (-2, 1) and (5, -5) and you have to find answers to a,b,c, and d using them two points

OpenStudy (anonymous):

Okay Hold On ,

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

m=y2−y1x2−x1→−5−15−(−2)→−5−15+2 What's the slope?

OpenStudy (anonymous):

is that for a

OpenStudy (anonymous):

Yeaa ,

OpenStudy (anonymous):

SLOPE=(-5-1)/(5+2)=-6/7